Use the power series method to find the general solution near x = 0 of (x2 + 4)y + xy = x + 2.
Answer #1

-(A-2) 2 (n=4)? .... ze az Let y = {anan be a power series solu to (2²+4)y+xy=x+2 near X0 mao Then y ni Enana in the given d